The following results represent the oxygen binding activity of purified myoglobin,

purified hemoglobin, and hemoglobin in human blood cells.

pO2 (Torr)

Fraction of purified myoglobin with O2

pO2 (Torr)

Fraction of purified hemoglobin with O2

pO2 (Torr)

Fraction of hemoglobin in red blood cells with O2

0.5

0.161

0.1

.00315

10.6

0.10

1.0

0.277

0.35

.0099

19.5

0.30

2.0

0.434

0.794

.0306

27.4

0.50

3.0

0.535

1.748

.0909

37.5

0.70

4.0

0.605

2.884

.24

50.4

0.85

6.0

0.697

4.467

.50

77.3

0.96

8.0

0.754

5.75

.76

92.3

0.98

12

0.821

7.94

.909

20

0.885

12.88

.969

29.51

.99

67.60

.997

Based on the data shown in the Table, determine the Hill coefficients of purified myoglobin, purified hemoglobin, and hemoglobin in red blood cells.
Answer #1

Answer:

  • Equation for Hill coficient = Y/1-Y = (pO2/P50)n
  • Taking the log of both side : log(Y/1-Y)= n(log pO2)- n(log P50)
  • Where Y = Fractional saturation of heamoglobin & 1-Y = free Heamoglobin

For purified Mayoglobin the Hill cofficient is-

  • n (pO2/P50) = n(log pO2)-  n(log P50) , here n=1
  • 1(log 20)- 1(log 2) = 1(1.30) - 1(0.30) =1 Hill cofficient n =1

For purified Heamoglobin the Hill cofficient is-

  • n log(pO2/P50)=  n(log pO2)-  n(log P50) , here n=4
  • 4(log 67.60)- 4(log 4.46) = 4(1.82) - 4(0.64)
  • 7.28 - 2.56 = 4.72 Hill cofficient n =4.72

For purified Heamoglobin in blood the Hill cofficient is-

  • n log(pO2/P50)=  n(log pO2)-  n(log P50) , here n=4
  • 4(log 92.3)- 4(log 27.4) = 4(1.69) - 4(1.43)
  • 6.76 - 5.72 =1.04 Hill cofficient n =1.04
  • n=1 describes no cooperative

n>1 describes positive cooperativity ( 1 ligand binding faciliates other)

n<1 describes negative cooperativity (1 ligand binding inhibits other binding)

Please Rate My Answer....... Thank......u...